# Syntax-Guided Procedural Synthesis of Molecules

### v1

@@ E.1 Featurization
$\text{BB}(n)=\text{FP}_{2048}(n_{\text{SMILES}})$ if $n$ is attributed with a building block from $\mathcal{B}$ or $\textbf{0}_{2048}$ otherwise and $\text{RXN}(n)=\text{one\_hot}_{91}(n_{\text{RXN\_ID}})$ if $n$ is attributed with a reaction from $\mathcal{R}$ or $\textbf{0}_{91}$ otherwise. We featurize each building block in $\mathcal{B}$ with their 256-dimension Morgan Fingerprint.
If $\mathcal{N}(T^{\prime})$ and $\mathcal{E}(T^{\prime})$ denote the node and edge set of $T^{\prime}\in\mathcal{T^{\prime}}$, then we define, for convenience:
$RXN({T^{\prime}})$ :=$:=\{r\in\mathcal{N}(T^{\prime})\mid\exists c,p\in\mathcal{N}(T^{\prime})\text{ s.t. }(r,c)\in\mathcal{E}(T^{\prime})\cap(p,r)\in\mathcal{E}(T^{\prime})\}$ (1)
$BB({T^{\prime}})$ =$=\{b\in\mathcal{N}(T^{\prime})\mid\nexists c\text{ s.t. }(b,c)\in\mathcal{E}(T^{\prime})\}.\}$ (2)
E.2 Loss Function
Each sample $(T^{\prime(i)},X^{(i)},y^{(i)})\in D_{\text{pretrain}}$. The loss can be specified as follows:
$\mathcal{L}_{D_{\text{pretrain}}}(\Phi)$ $\coloneqq\frac{1}{|D_{\text{pretrain}}|}\sum_{i=1}^{|D_{\text{pretrain}}|}\sum_{n\in RXN(T^{\prime(i)})}\text{CE}(F_{\Phi}(T^{\prime(i)})_{n},y^{(i)}_{n}),$
$\mathcal{L}_{D_{\text{pretrain}}}(\Omega)$ $\coloneqq\frac{1}{|D_{\text{pretrain}}|}\sum_{i=1}^{|D_{\text{pretrain}}|}\sum_{n\in BB(T^{\prime(i)})}\text{MSE}(F_{\Omega}(T^{\prime(i)})_{n},y^{(i)}_{n}).$
CE and MSE denote the standard cross entropy loss and mean squared error loss, respectively.
For evaluation, we define the following accuracy metrics:
$RXN_{\text{acc}}(\Phi)$ $=\frac{1}{|D_{\text{pretrain}}|}\sum_{i=1}^{|D_{\text{pretrain}}|}\frac{1}{|RXN(T^{\prime(i)})|}\sum_{n\in RXN(T^{\prime(i)})}\mathds{1}[\operatorname*{arg\,max}(F_{\Phi}(T^{\prime(i)})_{n})==\operatorname*{arg\,max}(y^{(i)}_{n})],$
$NN_{\text{acc}}(\Omega)$ $=\frac{1}{|D_{\text{pretrain}}|}\sum_{i=1}^{|D_{\text{pretrain}}|}\frac{1}{|NN(T^{\prime(i)})|}\sum_{n\in NN(T^{\prime(i)})}\mathds{1}[\operatorname*{arg\,min}_{B\in\mathcal{B}}\text{dist}(F_{\Omega}(T^{\prime(i)})_{n},\text{FP}_{256}(B))==n_{\text{SMILES}}]$
where we use the cosine distance.
E.3 Auxiliary Training Task
In 3.3.1, we defined the representation $T^{\prime}$ to be the parse tree of a partial program. However, we omitted an extra step that was used to preprocess $T^{\prime}$ for training.
To motivate this step, we first note the distinction between our program synthesis formulation and other formulations. Retrosynthesis is essentially already guided by the execution state at every step. Each expansion in the search tree executes a deterministic reaction template to obtain the new intermediate molecule. Planners based on single-step models [7], for example, assume the Markov Property by training models to directly predict a template given only the intermediate [63, 64]. In program synthesis, meanwhile, the state space is a set of partial programs with actions corresponding to growing the program. The execution of the program (or verification against the specification) does not happen until a complete program is obtained. In recent years, neural program synthesis methods found using auxiliary information in the form of the execution state of a program can help indirectly inform the search [5, 8, 17] since it gives a sense on what the program can compute so far. This insight does not apply to retrosynthesis, since retrosynthesis already executes on the fly. It also does not apply for the methods introduced in 2.3 that construct a synthetic tree in a bottom-up manner, for the same reason (the only difference is they use forward reaction templates, with a much smaller set of robust reaction templates) to obtain the execution state each step. However, as described in 3.3.1, our approach combines the computational advantages of restricting to a small set of forward reaction templates with the inductive bias of retrosynthetic analysis. Our policy is to predict forward reaction templates in a top-down manner. This formulation is common in top-down program synthesis, where an action corresponds to selecting a hole in the program. Similarly, our execution of the program does not happen until the tree is filled in. However, we leverage the insight that the execution state helps in an innovative way. We add an additional step when preprocessing $D_{\text{pretrain}}$: For each $T^{\prime}$ in $D_{\text{pretrain}}$, for each node $r$ corresponding to a reaction, we add a new node $o_{r}$ corresponding to the intermediate outcome of the reaction. If $RXN_{T^{\prime}}$ is the reaction nodes of $T^{\prime}$ as defined in E.1, we can construct $T^{\prime\prime}$ from $T^{\prime}$ as follows:
$\mathcal{N}(T^{\prime\prime})$ $\leftarrow\mathcal{N}(T^{\prime})\cup RXN_{T^{\prime}}$ (3)
$\mathcal{E}(T^{\prime\prime})$ $\leftarrow\mathcal{E}(T^{\prime})\cup\{(\text{parent}(r),o_{r}),(o_{r},r)\forall r\in RXN_{T^{\prime}}\}$ (4)
Figure 11: Examples of $T^{\prime\prime}$ where prediction targets are the frontier reactions (yellow circles), frontier building blocks (numbered yellow squares) and auxiliary intermediates (un-numbered yellow squares).
Lastly, we attribute each $o_{r}$ with the intermediate obtained from the original synthetic tree, i.e. executing the output of the program rooted at $r$. We featurize $\{y_{o}:=\text{FP}_{256}({o}_{\text{SMILES}})\}$ and add them as additional prediction targets to $D_{\text{pretrain}}$. Examples of $T^{\prime\prime}$ are given in 11.
E.4 Ablation Study
To understand whether the two key design choices for $\mathcal{T^{\prime\prime}}$ are justified, we did two ablations:
  1. 1.
    We use the original description of $\mathcal{T^{\prime}}$ in 3.3.1, i.e. without the auxiliary task.
  2. 2.
    We use $\mathcal{T^{\prime\prime}}$, but without attributing the intermediate nodes (so the set of targets is the same as Ablation 1.)
(a) Examples from $\mathcal{T^{\prime\prime}}$
(b) NN accuracy loss over top example 12(a)
(c) NN accuracy over bottom example 12(a)
(d) NN accuracy over $D_{\text{valid}}$
As shown in 12(d), using $\mathcal{T^{\prime\prime}}$ (Ours) achieves higher NN accuracy. This shows the benefit of learning the auxiliary training task. Meanwhile, ablating the auxiliary task (-aux) and ablating the intermediate node (-interm) does not have meaningful difference, indicating our architecture is robust to graph edits which are semantically equivalent. To understand the comparative advantage vs disadvantage of the auxiliary training task, consider the two examples in 12(a). The first example is equivalent to learning a single-step backward reaction prediction on forward templates [^1] For some templates, the forward template is one-to-one. For others, applying the backward template results in an ill-defined precursor, due to the many-to-one characteristic of these templates.. Our model clearly benefits from the auxiliary training task, which provides additional examples for learning the backward reaction steps. However, our model fares worse on predicting the first reactant of the top reaction. This may be due to competing resources. Despite the task being the same (and the set of forward templates are fixed), the model has to allocate sufficient capacity for the auxiliary task, whose output domain is much higher dimensional than $\mathcal{B}$. Ensuring positive transfer from learning the auxiliary task is an interesting extension for future work.
E.5 Model Architecture
@@
We opt for two Graph Neural Networks (for $\Phi,\Omega$), each with 5 modules. Each module uses a TransformerConv layer [54] (we use 8 attention heads), a ReLU activation, and a Dropout layer. We adopt sinusoidal positional embeddings via numbering nodes using the postorder traversal (to preserve the pairwise node relationships for all instances of the same skeleton). Then, we pretrain $\Phi,\Omega$ with $D_{\text{pretrain}}$.
E.6 Attention Visualization
@@
    If decoding a reaction, the reaction node which $F_{\Phi}$ predicts with the highest probability is decoded.
@@
    If decoding a building block, the node where the embedding from $F_{\Omega}$ has minimal distance to a building block is decoded.
In [54], each TransformerConv layer $l$ produces an attention weight for each edge, $[\alpha^{(l)}_{i,j}]$ where $\sum_{j}\alpha^{(l)}_{i,j}=1$. We average over all layers to obtain the mean attention weight for each directed edge, i.e., we set the thickness of each edge $(i,j)$ in each subfigure of 13 to be proportional $\sum_{l}\alpha^{(l)}_{i,j}$.
Figure 13: Case Study of Attention Flow
@@
    In 13(b), we see that $8$ attends significantly to the target, unsurprisingly. $8$ also attends to both its children, and attends more to its left child, which is a prior consistent with our general observation.
@@
    In 13(c), we see that after a specific reaction is instantiated at $8$, the attention dynamics somewhat change. The edge from $8$ to its left child thickens, while the edge from the left child to $8$ thins. This is likely because now that the identity of $8$ is known, it no longer needs to attend to its left child. The reciprocal relationship now intensifies, as the first reactant of $8$ now attends to $8$.
@@
    In 13(d), after the reaction at $6$ is decoded, we see the information propagate back up the tree and to the other subtree to inform $2$. We see the edge along the path from $6-8$ thickens, indicating the representation of $8$ is informed with new information, and in turn propagates it to $2$.
@@
    In 13(e), after the reaction at $2$ is decoded, we see the same phenomenon happen, where the information flow again propagates back up and to the other subtree. However, we see this comes with a tradeoff, as $6$ attends to its parent less, and instead reverts to its original attention strength to its children. We hypothesize the identity of $2$ has a strong effect on the posterior of $6$. This is an example where branching out to try more possible orders of decoding would facilitate a more complete algorithm.
@@
    In 13(f), we see how determining $5$ causes $6$ to attend more to $5$ than it does to its parent. Knowledge of $5$ allows the explaining away of $4$.
@@
    In 13(g), we note instances of a general phenomenon: the second reactant is decoded followed by the first. Empirically, the distribution of the second reactant has lower entropy than the first. $4$ was inferred after $5$ as the knowledge of its parent reaction and sibling reactant likely constrains its posterior significantly.
@@
    In 13(h), we see a similar phenomenon where the representation of $2$ attends slightly more to $1$ after it is decoded.
@@
Syntactic mutation Parent selection prob. of $i$ $\propto$ $(\mathrm{rank}(i)+10)^{2}$
Num. tree edits $n_{\text{edit}}$ $\mathcal{U}\{1,2,3\}$
@@
Appendix F Genetic Algorithm
Our genetic algorithm (GA) is designed to mimic SynNet’s [21], and settings are given in Table 5. We fix the same number of offspring fitness evaluations per generation to ensure a fair comparison, strategically allocating the evaluations between offspring generated using Semantic Evolution and those generated using Syntactic Mutation.
F.1 Semantic Evolution
Given two parents $(X_{1},T_{1})$ and $(X_{2},T_{2})$, semantic evolution samples a child $(X,T)$ as follows. We obtain $X$ by combining $n_{\text{cross}}$ random bits from $X_{1}$ and the other $2048-n_{\text{cross}}$ bits from $X_{2}$ and then, with probability $p_{\text{flip}}$, flipping $n_{\text{flip}}$ random bits of the crossover result. We set $T=Q_{\Theta}(X)$ using the recognition model.
F.2 Syntactic Mutation
Given an individual $(X,T)$ from F.1, syntactic mutation mutates $T$ to obtain a syntactic analog. We perform $n_{\text{edit}}$ edits on $T$ to obtain $T^{\prime}$. With equal probability, each edit either adds or removes a random leaf. To do so, we enumerate all possible additions and removals, and ignore the ones that produce a mutant tree with less than 2 nodes or more than 4 internal nodes. The edit is uniformly sampled from all such choices, or no operation is performed if no viable choices exist. In the early iterations of the GA, we set $T\leftarrow T^{\prime}$ using the criteria that $\text{exec}(F_{\Phi,\Omega}(X,T^{\prime}))$ promotes a higher Internal Diversity than $\text{exec}(F_{\Phi,\Omega}(X,T))$. In the later iterations of the GA, we use the collected experience of iterations prior to fit a surrogate Gaussian Process. We set $T\leftarrow T^{\prime}$ and accept the mutant if $\text{exec}(F_{\Phi,\Omega}(X,T^{\prime}))$ has higher acquisition value than a counterpart obtained through Semantic Evolution. We refer the reader to work on Particle Swarm Optimization with Gaussian Processes [62, 71, 30] for the related literature.
F.3 Surrogate Checkpoint
The surrogate checkpoint was trained using $D_{\text{pretrain}}$ as described in Appendix E. To lower the runtime of the GA, we only reconstruct using a random subset of the input skeleton’s possible topological orders. For each topological order, we follow a greedy decoding scheme where reactions are decoded before building blocks, as described in Appendix E.6.
Appendix G Full Results on TDC Oracles
@@
Table 9: Guacamol multi-objective Oracles for properties of known drugs: Perindopril, Amlodipine, Sitagliptin, and Zaleplon.
Tables 6, 7, 8 and 9 are comprehensive results against baselines taxonomized in [22]. We evaluate the average score of the Top K molecules, their average synthetic accessibility [18] and top K AUC (AUC of no. oracle calls vs score plot), for K=1,10,100. Like [22], we limit to 10000 Oracle calls, truncating and padding to 10000 if convergence occurs before 10000 calls. For each cell, numbers are followed by rankings. $X(R_{1}|R_{2})$ means score $X$ is ranked $R_{1}$-best amongst all methods for that column and $R_{2}$-best amongst in-category methods. We visualize the rankings in Figure 14 to facilitate easier interpretation of the results.
@@
Appendix H Docking Case Study with AutoDock Vina

### v2

@@ F.5.2 Results on Synthesizable Molecular Design
Top 1 Top 10 Top 100 Top 1 Top 10 Top 100
category Oracle Calls Score SA AUC Score SA AUC Score SA AUC Oracle Calls Score SA AUC Score SA AUC Score SA AUC
Ours (MC) synthesis 5056 0.98 2.045 0.923 0.97 2.294 0.893 0.942 2.294 0.814 7949 0.4 4.12 0.356 0.342 3.902 0.304 0.295 4.013 0.256
Ours (MC10) synthesis 8045 0.4 3.353 0.357 0.344 4.593 0.297 0.302 4.44 0.247
Ours synthesis 4886 0.98 2.045 0.891 0.967 2.302 0.848 0.944 2.27 0.778 8303 0.4 3.353 0.371 0.342 4.161 0.305 0.298 4.256 0.252
Osimertinib MPO Perindopril MPO
Top 1 Top 10 Top 100 Top 1 Top 10 Top 100
category Oracle Calls Score SA AUC Score SA AUC Score SA AUC Oracle Calls Score SA AUC Score SA AUC Score SA AUC
Ours (MC) synthesis 9402 0.865 2.282 0.830 0.853 2.187 0.813 0.841 2.189 0.771 10000 0.572 3.101 0.541 0.567 3.072 0.521 0.555 3.077 0.486
Ours (MC10) synthesis 5056 0.98 2.045 0.923 0.97 2.294 0.893 0.942 2.294 0.814 10000 0.596 3.263 0.542 0.574 3.147 0.523 0.556 3.058 0.489
Ours synthesis 10000 0.859 2.263 0.826 0.847 2.21 0.81 0.832 2.249 0.769 10000 0.622 3.338 0.547 0.591 3.378 0.524 0.558 3.137 0.485
Depends on the task. We also include preliminary results on synthesizable molecular design. We find the results hold up for the MC models. Encouragingly, for the difficult task of Osimertinib MPO, we see for $C=10$, the results are substantially better. At the same time, Ours remain better for Perindopril MPO, which suggests each training strategy may suit different tasks. We also see for easier tasks like GSK and Median 1, the $C=1$ model and Ours are essentially interchangeable. This suggests even a low reconstruction accuracy suffices for these Oracles.
Implications. The stratified sampling is designed to enhance supervised policy learning on larger programs while preventing the combinatorial explosion in Algo. 1 and to some extent, overfitting. However, this may benefit easier tasks (e.g. distributions of synthesizable molecules) while disadvantaging harder tasks. Thus, $C$ should be tuned for optimal trade-off between efficiency and downstream synthesizable analog generation performance.
F.6 Model Architecture
@@
We opt for two Graph Neural Networks (for $\Phi,\Omega$), each with 5 modules. Each module uses a TransformerConv layer (Shi et al., 2020) (we use 8 attention heads), a ReLU activation, and a Dropout layer. We adopt sinusoidal positional embeddings via numbering nodes using the postorder traversal (to preserve the pairwise node relationships for the same skeleton). Then, we pretrain $\Phi,\Omega$ with ${\mathcal{D}}$.
F.7 Training & Convergence Analysis
(a) Legend
(b) NN training step vs NN accuracy
(c) RXN training step vs accuracy
Figure 15: We plot the number of training steps needed to converge our models under the standard and linear (-MC) training strategies for $C=1$.
We elaborate on 4.1.3 further with a quantitative comparison of training costs with SynNet Gao et al. (2022). The key difference is $\mathcal{D}_{\text{synnet}}$ batches by the size of a synthetic tree, whereas we batch by the synthetic trees (in program form).
SynNet Training Complexity: SynNet serializes the construction of a synthetic tree, so a training epoch does $O(\sum_{\mathcal{D}_{0}}|T|)$ passes, where $|T|$ is the number of nodes (or number of edges, as they are different by one). The cost of a MLP forward and backward pass for SynNet is $O(L\cdot H^{2}+F\cdot H)$. The total complexity per epoch is $O((\sum|T|)(L\cdot H^{2}+F\cdot H))$.
SynthesisNet Linear Training Complexity: As discussed in 4.1.3, a training epoch does $O(\mathcal{D}_{0})$ passes, where the constant factor can be adjusted. The cost of a GNN forward and backward pass on a tree $T$ is $O(L\cdot|T|\cdot(F\cdot H+H^{2}))$, where $L$, $F$, $H$ are number of layers, feature and hidden dimension, respectively. The total complexity per epoch is $O(L\cdot(\sum|T|)\cdot(F\cdot H+H^{2}))$ since we train on the forest of trees over $\mathcal{D}_{0}$. This is equivalent to $O((\sum|T|)(L\cdot H^{2}+F\cdot H))$.
We can conclude the per-epoch complexity following the linear training strategy is equivalent to that of SynNet’s. However, what matters in practice is the convergence rate, so we also include a quantitative comparison between convergence plots. We find that even SynthesisNet’s standard training strategy is practically equivalent to SynNet in the number of passes needed to converge the model. As a disclaimer, we only include SynthesisNet’s numbers using the default setting of $k=4$, where $|\mathcal{D}_{0}|\approx 135k$ and $|\mathcal{D}|\approx 818k$. We also show results of linear training with constant factor of $1$ (i.e. $|\mathcal{D}|=|\mathcal{D}_{0}|$).
Convergence Comparison: Both SynNet and SynthesisNet uses a batch size of 64. We see from Figure 15 that SynthesisNet requires $\approx 170k$ training steps (batches) to converge (combining both BB and RXN networks) following standard training. Meanwhile, we refer readers to Figure 13 of Gao et al. (2021), where the convergence plots show at least (with the most generous interpretation) 1M steps needed to converge each of the action, reactant1, reaction and reactant2 networks. Pooling the training of all networks, we give a very generous estimate of $1M$ batches for SynNet to converge. What’s left is to figure the average scaling size factor from a batch of trees (ours) to a batch of tree nodes (SynNet), which computes as $818k/135k\approx 6$, which implies $\approx 6\cdot 170k\approx 1M$ steps. We can conclude the SynthesisNet with standard training is comparable if not more efficient in the number of training steps to converge the model. Should efficiency be of further concern, we suggest using the Linear Training strategy, and adjusting the sample constant factor accordingly.
Training and Inference Time: Converging the RXN and BB networks took us $\approx$ 1 and 5 hours on a single NVIDIA RTX A6000. A single inference call to the surrogate takes a few seconds.
F.8 Attention Visualization
@@
    If decoding a reaction, the reaction node which $\pi_{\mathcal{R}}$ predicts with the highest probability is decoded.
@@
    If decoding a building block, the node where the embedding from $\pi_{\mathcal{B}}$ has minimal distance to a building block is decoded.
In (Shi et al., 2020), each TransformerConv layer $l$ produces an attention weight for each edge, $[\alpha^{(l)}_{i,j}]$ where $\sum_{j}\alpha^{(l)}_{i,j}=1$. We average over all layers to obtain the mean attention weight for each directed edge, i.e., we set the thickness of each edge $(i,j)$ in each subfigure of Figure 16 to be proportional $\sum_{l}\alpha^{(l)}_{i,j}$.
@@
    In Figure 16(b), we see that $8$ attends significantly to the target, unsurprisingly. $8$ also attends to both its children, and attends more to its left child, which is a prior consistent with our general observation.
@@
    In Figure 16(c), we see that after a specific reaction is instantiated at $8$, the attention dynamics somewhat change. The edge from $8$ to its left child thickens, while the edge from the left child to $8$ thins. This is likely because now that the identity of $8$ is known, it no longer needs to attend to its left child. The reciprocal relationship now intensifies, as the first reactant of $8$ now attends to $8$.
@@
    In Figure 16(d), after the reaction at $6$ is decoded, we see the information propagate back up the tree and to the other subtree to inform $2$. We see the edge along the path from $6-8$ thickens, indicating the representation of $8$ is informed with new information, and in turn propagates it to $2$.
@@
    In Figure 16(e), after the reaction at $2$ is decoded, we see the same phenomenon happen, where the information flow again propagates back up and to the other subtree. However, we see this comes with a tradeoff, as $6$ attends to its parent less, and instead reverts to its original attention strength to its children. We hypothesize the identity of $2$ has a strong effect on the posterior of $6$. This is an example where branching out to try more possible orders of decoding would facilitate a more complete algorithm.
@@
    In Figure 16(f), we see how determining $5$ causes $6$ to attend more to $5$ than it does to its parent. Knowledge of $5$ allows the explaining away of $4$.
@@
    In Figure 16(g), we note instances of a general phenomenon: the second reactant is decoded followed by the first. Empirically, the distribution of the second reactant has lower entropy than the first. $4$ was inferred after $5$ as the knowledge of its parent reaction and sibling reactant likely constrains its posterior significantly.
@@
    In Figure 16(h), we see a similar phenomenon where the representation of $2$ attends slightly more to $1$ after it is decoded.
@@
Syntactic mutation Num. tree edits $n_{\text{edit}}$ $\mathcal{U}\{1,2,3\}$
@@
Appendix G Genetic Algorithm
Our genetic algorithm (GA) is designed to mimic SynNet’s (Gao et al., 2021), and settings are given in Table 7. We fix the same number of offspring fitness evaluations per generation to ensure a fair comparison, strategically allocating the evaluations between offspring generated using semantic evolution and those generated using syntactic mutation.
G.1 Semantic Evolution
Given two parents ${\bm{x}}_{1}$ and ${\bm{x}}_{2}$, semantic evolution samples a child ${\bm{x}}_{*}$ as follows. We combine $n_{\text{cross}}$ random bits from ${\bm{x}}_{1}$ and the other $2048-n_{\text{cross}}$ bits from ${\bm{x}}_{2}$ and then, with probability $p_{\text{flip}}$, flipping $n_{\text{flip}}$ random bits of the crossover result.
G.2 Syntactic Mutation
Given a child ${\bm{x}}_{*}$ from Appendix G.1, syntactic mutation performs $n_{\text{edit}}$ edits on $T=\tau{}({\bm{x}}_{*})$ to obtain a syntactic analog $T$. With equal probability, each edit either adds or removes a random leaf. To do so, we enumerate all possible additions and removals, and ignore the ones that produce an empty tree or a tree with more than 4 reaction nodes. The edit is uniformly sampled from all such choices, or no operation is performed if no viable choices exist. Using the surrogate, the siblings $({\bm{x}}_{*},T)$ and $({\bm{x}}_{*},T)$ are then turned into two fingerprints, and one of with the higher expected improvement under a Gaussian process (GP) is selected. Our GP uses a radial basis function kernel with length scale 1 and is fitted using the population and offspring from the preceding generation.
G.3 Surrogate Checkpoint
The surrogate checkpoint was trained as described in Appendix F. To lower the runtime of the GA, we only reconstruct using a random subset of the input skeleton’s possible topological orders. For each topological order, we follow a greedy decoding scheme where reactions are decoded before building blocks, as described in Appendix F.8.
Appendix H Full Results on TDC Oracles
@@
Table 11: Guacamol multi-objective Oracles for properties of known drugs: Perindopril, Amlodipine, Sitagliptin, and Zaleplon. Disclaimer: We used version 1.0.0. PMO Gao et al. (2022), where we adopted the numbers from, used 0.3.6. We’d like to mention there were changes since 0.3.6 that may affect the Sitagliptin and Zaleplon evaluations.
@@
Tables 8, 9, 10 and 11 are comprehensive results against baselines taxonomized in (Gao et al., 2022). We evaluate the average score of the Top K molecules, their average synthetic accessibility (Ertl & Schuffenhauer, 2009) and top K AUC (AUC of no. oracle calls vs score plot), for K=1,10,100. Like (Gao et al., 2022), we limit to 10000 Oracle calls, truncating and padding to 10000 if convergence occurs before 10000 calls. For each cell, numbers are followed by rankings. $X(R_{1}|R_{2})$ means score $X$ is ranked $R_{1}$-best amongst all methods for that column and $R_{2}$-best amongst in-category methods. We visualize the rankings in Figure 17 to facilitate easier interpretation of the results.
@@
Appendix I Docking Case Study with AutoDock Vina